OpenStudy (anonymous):

If a person draws 2 cards at random from a standard deck of 52 cards, what is the probability that they get two aces?

OpenStudy (anonymous):

is this correct 4C2 / 52C2

OpenStudy (anonymous):

@hartnn @satellite73

OpenStudy (anonymous):

They will have a 4 in 52 chance to draw an ace on the first card, followed by a 3 in 51 chance on the second: \[\frac{4}{52}\frac{3}{51}=\frac{12}{2652}=0.004524\] your choose notation gives the same results

OpenStudy (anonymous):

So...My answer is correct @Mathmuse

OpenStudy (anonymous):

but 4C2 = 4!/2! * 2! = 3 52C2 = 52! / 50! 2! = 26 * 51 3/ 26 * 51 = 3/26 * 17 = 1/442

OpenStudy (anonymous):

@Mathmuse @Mathmuse @Mathmuse @Mathmuse @Mathmuse @Mathmuse @Mathmuse @Mathmuse @Mathmuse @Mathmuse @Mathmuse @Mathmuse @Mathmuse @Mathmuse

OpenStudy (anonymous):

\[4C2=\frac{4!}{2!(4-2)!}=\frac{24}{2(2)}=6\]